OpenStudy (anonymous):

slove: 3x-4y=10 5x-3y=2

OpenStudy (turingtest):

3x-4y=10 (1) 5x-3y=2 (2) multiply (1) by -3 and (2) by 4: -9x+12y=-30 10x-12y=80 add (1) to (2): x=50 solve either equation for y: 3x-4y=10 y=(3/4)x-5/2=3/4(50)-5/2=70/2=35 so x=50 y=35 is the solution

OpenStudy (anonymous):

thers a slight error on 5th line turing it 8 not 80

OpenStudy (turingtest):

Oh man that's not slight. Dang...